# TS-PEFT: Token-Selective Parameter-Efficient Fine-Tuning with Learnable Threshold Gating

### v1

@@ Abstract
In the field of large models (LMs) for natural language processing (NLP) and computer vision (CV), Parameter-Efficient Fine-Tuning (PEFT) has emerged as a resource-efficient method that modifies a limited number of parameters while keeping the pretrained weights fixed. This paper investigates the traditional PEFT approach, which applies modifications to all position indices, and questions its necessity. We introduce a new paradigm called Token-Selective PEFT (TS-PEFT), in which a function S selectively applies PEFT modifications to a subset of position indices, potentially enhancing performance on downstream tasks. Our experimental results reveal that the indiscriminate application of PEFT to all indices is not only superfluous, but may also be counterproductive. This study offers a fresh perspective on PEFT, advocating for a more targeted approach to modifications and providing a framework for future research to optimize the fine-tuning process for large models.
@@ 1 Introduction
Large models (LMs) have achieved significant progress in natural language processing (NLP) yao2024survey, computer vision (CV) croitoru2023diffusion; liu2024sora, and many other fields. However, due to their large parameter scales, directly fine-tuning all parameters is often impractical in terms of computational and storage overhead. To address this issue, parameter-efficient fine-tuning (PEFT) has gradually become a mainstream paradigm for adapting pretrained large models: by freezing the base model network and introducing only a small number of trainable parameters, downstream customization can be accomplished efficiently hu2023llm; hanparameter.
Various PEFT methods have demonstrated strong performance across a wide range of downstream tasks. Among them, prompt-based approaches (such as Prefix-Tuning li2021prefix and Prompt-Tuning lester2021power) adjust model behavior by learning additional input tokens, while low-rank adaptation (LoRA) hu2022lora and its variants (e.g., AdaLoRA zhang2023adaptive, DoRA liu2024dora, VeRA kopiczkovera) update internal representations by injecting a small number of trainable parameters or low-rank components into the base model. These approaches significantly reduce the number of trainable parameters while typically maintaining performance comparable to, or even approaching, full-parameter fine-tuning.
Most existing PEFT methods, however, implicitly rely on a key assumption: once a layer is selected as the target module for PEFT, all token positions passing through this module should receive a trainable modification. Formally, for an input sequence $X=\{x_{i}\}_{i=1}^{T}$ and a frozen weight matrix $W_{0}$, the hidden state at position $i$ is typically written as:
@@
where $\operatorname{M}(\cdot)$ is a PEFT module such as LoRA. This “update-all-positions” design ignores the fact that different token positions contribute unequally to the downstream objective.
@@
In practice, we observe that, for already strong pretrained models, the deviation induced by PEFT is extremely small on many token positions, while only a small subset of positions exhibit substantial and task-relevant changes. In other words, most PEFT updates are effectively redundant at the token level. Uniformly applying modifications to all positions not only injects unnecessary noise and potential interference, but also increases computation, especially when multiple PEFT modules are stacked. This motivates a natural question: can PEFT be made selective over token positions instead of blindly updating every one of them?
Our empirical findings suggest a negative answer. We propose a simple yet effective framework, Token-Selective PEFT (TS-PEFT), which enables token-level selective parameter-efficient fine-tuning. Instead of always applying $\operatorname{M}(x_{i})$ to every position, TS-PEFT introduces a binary gating function $S$ at each token position that decides whether to add the PEFT update or to keep the base model output unchanged. Intuitively, positions whose PEFT updates are small relative to the frozen output remain untouched, while positions with larger, more impactful updates retain their PEFT modifications, yielding a token-level sparse pattern of PEFT updates.
In the experimental section, we conduct systematic evaluations on representative models including LLaMA-3-8B, LLaVA-1.5-7B, and DeBERTaV3-base hedebertav3, covering commonsense reasoning, visual instruction tuning, and natural language understanding. The results show that, on top of strong PEFT baselines such as LoRA, DoRA, AdaLoRA, and Adapter, TS-PEFT consistently matches or surpasses their performance while applying updates to only about 40–60% of token positions. Moreover, the learned token-level sparsity patterns can be aggregated into module-level importance indicators, which enable selecting fewer yet more critical tuning modules that even outperform full-module fine-tuning under comparable parameter budgets. In summary, our contributions are as follows:
  • We point out and examine a neglected aspect of standard PEFT: once a layer is chosen as a target module, all token positions are updated uniformly, even though only a small fraction of positions contribute meaningfully to the downstream objective. Our empirical analysis of token-level sparsity shows that many PEFT updates induce negligible deviation from the frozen backbone, revealing substantial redundancy along the sequence dimension and motivating selective updates.
  • We propose TS-PEFT, which introduces a token-level binary gating function on top of existing PEFT modules. By formulating the learning of layer-wise thresholds as a sparsity-regularized proximal optimization problem, deriving a practical approximate gradient, and applying Adam-style moment estimation, TS-PEFT enables stable end-to-end training of token-level gates jointly with PEFT parameters. The framework is agnostic to the specific low-rank parameterization and can be used as a plug-in module for a variety of PEFT methods.
  • On three representative benchmarks, TS-PEFT consistently matches or surpasses strong PEFT baselines, while updating only 40%-60% token positions. Moreover, the learned token-level sparsity patterns can be aggregated into module-wise importance scores, enabling selective tuning of fewer but more critical modules that can outperform full-module fine-tuning under comparable parameter budgets.
@@ 2 Related Work
Low-Rank Adaptation. Low-rank adaptation and its derivatives have demonstrated strong potential in reducing the number of trainable parameters while preserving downstream performance. LoRA hu2022lora introduces two low-rank trainable matrices $A$ and $B$ to approximate weight updates, thereby avoiding direct modification of full weight matrices in large pretrained models; after training, these matrices can be merged back into the base model without incurring additional inference cost. Building on this idea, AdaLoRA zhang2023adaptive adaptively redistributes the effective rank across different layers to better meet task-specific requirements, DoRA liu2024dora refines the decomposition of weight updates to improve the expressiveness and stability of low-rank adaptation, and VeRA kopiczkovera leverages shared low-rank bases to further reduce per-task parameter overhead. However, all these methods implicitly assume that once a layer is selected as the PEFT target, every token position passing through that layer should receive an update—that is, they perform parameter compression at the module level while still maintaining dense updates along the positional dimension. This may introduce unnecessary redundant modifications and fails to explicitly characterize or exploit potential token-level redundancy.
Sparsity. Sparsity-guided methods have been widely used to improve the efficiency and interpretability of large language models. By reducing the number of active parameters and concentrating computation on core components, these approaches not only make models more efficient but also offer clearer insights into their decision-making processes. Mixture-of-experts (MoE) models, exemplified by Switch Transformers fedus2022switch and Mixtral 8x7B jiang2024mixtral, employ sparse activation to keep the computational cost roughly constant even with a very large number of parameters, and several recent works have combined MoE architectures with LoRA-style adapters to further enhance performance chen2024llava; liu2024moe; zadouripushing . Another typical use of sparsity is model pruning, where less influential parameters are removed while preserving important ones to maintain accuracy under lower computational budgets ma2023llm; xubesa; xiasheared. tan2024sparsity further introduce sparsity-guided techniques as a tool for interpreting the behavior of LLMs. However, most existing studies focus on parameter-wise sparsity, i.e., deciding which weights or modules to activate or prune. In contrast, our work explores token-level sparsity in the context of PEFT: instead of sparsifying parameters, we selectively decide which token positions should receive PEFT updates, revealing and leveraging redundancy along the sequence dimension.
@@ 4.2 Experimental Results
TS-PEFT introduces token-level sparsity as a logical mask. Therefore, the current implementation does not provide wall-clock speedups. However, the theoretical computation reduction is significant. With specialized kernels support for sparse matrix additions in the future, TS-PEFT has the potential to translate this logical sparsity into physical latency reduction.
4.3 In-depth Analysis
Figure 2: Performance of LLaMA3.1-8B on CSR benchmarks as the percentage of selected modules varies for each PEFT method.
@@ 4.3 In-depth Analysis
Table 4: The performance of LLaMA3.1-8B on the commonsense reasoning task using various selection methods.
Necessity of Adaptive Moment Estimation. To verify the effect of the adaptive momentum estimation module, we conduct ablation experiments by removing this component. In this case, the threshold $\tau$ exhibits noticeable oscillation and gradually diverges, and the training loss also fails to converge. On representative CSR tasks such as BoolQ, this instability leads to a clear degradation in dev performance compared with the full TS-PEFT variant (see Appendix A.2 for detailed curves and numerical results). These observations indicate that adaptive momentum estimation is essential for stabilizing the threshold update process in practice.
Module-wise Sparsity Distribution. In LLaMA3.1-8B, there are 32 layers in total. We compute the sparsity of q_proj, k_proj, and v_proj in each layer and take their average to obtain the sparsity distribution of TS-PEFT across modules (detailed results are provided in Appendix A.3). Overall, the sparsity of TS-LoRA and TS-AdaLoRA is relatively evenly distributed across layers, with standard deviations of 9.8 and 11.4, respectively. In contrast, TS-DoRA exhibits a much higher standard deviation of 41.3, showing a more “polarized” pattern. It should be noted that LoRA and AdaLoRA share the same weight structure despite having different rank allocation strategies, whereas DoRA adopts a distinct parameterization by decomposing weights into magnitude and direction. We hypothesize that this structural difference leads to the stronger selectivity observed in TS-DoRA at the module level, naturally raising the question: are the modules with lower sparsity more critical for downstream tasks?
@@
Effect of the Number of Target Modules. After using sparsity as the criterion for module importance ranking, we further examine the effect of varying the proportion of target modules while keeping the ranking fixed: we select different percentages of modules in ascending order of sparsity and fine-tune them. The results are shown in Figure 2. For AdaLoRA, performance remains largely unaffected as the proportion of target modules increases. However, for the others, “more is better” does not hold: DoRA achieves its best performance when using only 50% of the modules, LoRA reaches its optimal performance at 80%, and even using only 20% of the modules yields results close to the optimum.
In addition, LoRA and DoRA exhibit clear performance cliffs at certain points: for DoRA, performance drops by about 6.6% when increasing the target module proportion from 50% to 60%; for LoRA, increasing the proportion from 90% to 100% results in a drop of about 1.3%. This phenomenon indicates that jointly training modules with higher sparsity may negatively interfere with overall performance. In contrast, AdaLoRA is less sensitive to such “low-importance modules,” as its dynamic rank allocation redistributes capacity across modules.
@@ 5 Conclusion
This paper revisits parameter-efficient fine-tuning from a token-level perspective and shows that standard PEFT methods exhibit substantial redundancy by uniformly updating all token positions within selected layers. We propose TS-PEFT, which introduces a learnable threshold gate to selectively apply PEFT updates to only about 40–60% of token positions, yet consistently matches or surpasses strong baselines. Our analyses further show that the resulting token-level sparsity patterns provide a meaningful signal of module importance, enabling more effective module selection under the same parameter budget and offering a basis for future work on more efficient sparse fine-tuning and inference.

### v2

# TS-PEFT: Unveiling Token-Level Redundancy in Parameter-Efficient Fine-Tuning

@@ Abstract
Current Parameter-Efficient Fine-Tuning (PEFT) methods typically operate under an implicit assumption: once a target module is selected, every token passing through it contributes equally to the downstream task and requires a parameter update. In this paper, we challenge this convention and unveil a pervasive token-level redundancy in the fine-tuning of large models. We propose TS-PEFT, a theoretically grounded framework utilizing proximal optimization to dynamically identify and skip redundant token updates during training. Our extensive experiments across Natural Language Understanding, Commonsense Reasoning, and Visual Instruction Tuning demonstrate that indiscriminately updating all tokens is not only computationally superfluous but often introduces optimization noise. Strikingly, by discarding 40%-60% of token updates, TS-PEFT consistently matches or surpasses the performance of dense baselines (e.g., LoRA, DoRA). Furthermore, we provide an in-depth analysis revealing that the learned token-level sparsity serves as a superior indicator of module importance compared to traditional weight norms, offering a novel data-driven perspective on the intrinsic adaptation mechanism of large models.
@@ 1 Introduction
Large models (LMs) have demonstrated remarkable capabilities across Natural Language Processing (NLP) yao2024survey and Computer Vision (CV) croitoru2023diffusion; liu2024sora. To adapt these massive models to downstream tasks efficiently, Parameter-Efficient Fine-Tuning (PEFT) has become the de facto paradigmhu2023llm; hanparameter. Among them, prompt-based approaches (such as Prefix-Tuning li2021prefix and Prompt-Tuning lester2021power) adjust model behavior by learning additional input tokens, while low-rank adaptation (LoRA) hu2022lora and its variants (e.g., AdaLoRA zhang2023adaptive, DoRA liu2024dora, VeRA kopiczkovera) update internal representations by injecting a small number of trainable parameters or low-rank components into the base model. These approaches significantly reduce the number of trainable parameters while typically maintaining performance comparable to, or even approaching, full-parameter fine-tuning.
However, despite their distinct architectural designs, these methods share a common, implicit assumption: the ”Dense Update” assumption. That is, once a module (e.g., a self-attention layer) is targeted for fine-tuning, the learnable modification is indiscriminately applied to every token position in the input sequence. Formally, for an input sequence $X=\{x_{i}\}_{i=1}^{T}$ and a frozen weight matrix $W_{0}$, the hidden state at position $i$ is typically written as:
@@
where $\operatorname{M}(\cdot)$ is a PEFT module such as LoRA.
@@
In this paper, we challenge this convention and investigate a fundamental scientific question: Is it necessary to apply fine-tuning updates to all token positions? Given the strong representational power of pre-trained backbones, we hypothesize that for a specific downstream sample, only a subset of tokens requires task-specific adaptation, while others retain sufficient information from the pre-trained weights. Under this hypothesis, the standard dense update strategy is not only computationally superfluous but may also introduce optimization noise and interference, potentially hindering model performance.
Recent works like CFT ruan2025enhancing have demonstrated that masking out non-critical tokens during training improves reasoning. However, CFT treats redundancy as a static data filtering problem, requiring costly offline counterfactual perturbations to generate fixed masks. In contrast, to systematically verify this hypothesis and quantify the potential redundancy, we propose Token-Selective PEFT (TS-PEFT). Unlike prior approaches, TS-PEFT serves as a learnable probe that dynamically decides whether to update a token or keep its representation frozen. We formulate this selection process as a sparsity-regularized optimization problem, employing proximal optimization and an Adam-style adaptive thresholding mechanism to ensure stable, end-to-end training of the gating decisions.
Our empirical findings unveil a pervasive token-level redundancy in standard PEFT. Across diverse benchmarks including Commonsense Reasoning, Visual Instruction Tuning, and Natural Language Understanding, we observe that discarding 40%-60% of token updates does not degrade performance. Strikingly, TS-PEFT often outperforms its dense counterparts (e.g., LoRA, DoRA) by explicitly filtering out redundant updates. While current hardware constraints on sparse operations limit immediate wall-clock speedups, our work establishes a theoretical computational bound, demonstrating that nearly half of the fine-tuning computation in current paradigms is effectively unnecessary.
Furthermore, our analysis exposes a profound connection between token sparsity and module importance. We discover that the learned sparsity pattern is not random; modules that exhibit low sparsity (i.e., require frequent updates) are significantly more critical to the downstream task than those with high sparsity. This insight allows token-level sparsity to serve as a superior indicator for module selection compared to traditional weight-norm-based metrics.
In summary, our contributions are as follows:
  • Unveiling Redundancy: We identify and quantify that nearly 40%-60% of token updates in standard PEFT are redundant. We empirically demonstrate that indiscriminately updating these tokens is suboptimal and introduces optimization noise.
  • Methodological Framework: We propose TS-PEFT, a theoretically grounded framework that employs proximal optimization to dynamically learn token-selective policies in an end-to-end manner, bypassing the need for costly offline heuristics.
  • Mechanism Insight: We reveal that token-level sparsity provides a strong signal for module importance, offering a novel data-driven perspective on the intrinsic adaptation mechanism of large models.
@@ 2 Related Work
Low-Rank Adaptation. Low-rank adaptation and its derivatives have demonstrated strong potential in reducing the number of trainable parameters while preserving downstream performance. LoRA hu2022lora introduces two low-rank trainable matrices $A$ and $B$ to approximate weight updates, thereby avoiding direct modification of full weight matrices in large pretrained models; after training, these matrices can be merged back into the base model without incurring additional inference cost. Building on this idea, AdaLoRA zhang2023adaptive adaptively redistributes the effective rank across different layers to better meet task-specific requirements, DoRA liu2024dora refines the decomposition of weight updates to improve the expressiveness and stability of low-rank adaptation, and VeRA kopiczkovera leverages shared low-rank bases to further reduce per-task parameter overhead. However, all these methods implicitly assume that once a layer is selected as the PEFT target, every token position passing through that layer should receive an update. That is to say, they perform parameter compression at the module level while still maintaining dense updates along the positional dimension. This may introduce unnecessary redundant modifications and fails to explicitly characterize or exploit potential token-level redundancy.
Sparsity. Sparsity-guided methods have been widely used to improve the efficiency and interpretability of large language models. By reducing the number of active parameters and concentrating computation on core components, these approaches not only make models more efficient but also offer clearer insights into their decision-making processes. Mixture-of-experts (MoE) models, exemplified by Switch Transformers fedus2022switch and Mixtral 8x7B jiang2024mixtral, employ sparse activation to keep the computational cost roughly constant even with a very large number of parameters, and several recent works have combined MoE architectures with LoRA-style adapters to further enhance performance chen2024llava; liu2024moe; zadouripushing . Another typical use of sparsity is model pruning, where less influential parameters are removed while preserving important ones to maintain accuracy under lower computational budgets ma2023llm; xubesa; xiasheared. tan2024sparsity further introduce sparsity-guided techniques as a tool for interpreting the behavior of LLMs.
We observe that these methods exploit redundancy at different levels. MoE-based approaches typically target modular-level sparsity, whereas pruning focuses on parameter-level sparsity. In contrast, our work explores token-level sparsity within the context of PEFT, introducing a gating mechanism to selectively skip updates for tokens requiring no task-specific adaptation. Therefore, our method is independent of these architectures and can be theoretically plugged in as a detachable module on top of LoRA-MoE, further eliminating redundant computations.
@@ 3.4 Adam-style Threshold Update and Overall Training Procedure
Optimization Stability. We empirically observe that the adaptive moment estimation (Eq. (16)-Eq. (20)) is critical for convergence. Ablation studies in Appendix A.2 demonstrate that removing this component leads to severe threshold oscillation and training divergence, resulting in significant performance degradation on sensitive tasks like BoolQ.
@@ 4.2 Experimental Results
Although TS-PEFT achieves a theoretical computation reduction of approximately 40%-60% in FLOPs within the adapter modules, current deep learning accelerators and libraries remain heavily optimized for dense matrix operations. Consequently, the imposed token-level sparsity currently serves as a theoretical computational bound and a probe for quantifying redundancy, rather than an immediate mechanism for wall-clock speedup. Nevertheless, with the emerging support for sparse tensor cores in next-generation hardware, TS-PEFT establishes a blueprint for translating this identified redundancy into tangible physical efficiency.
4.3 Analysis: Token Sparsity Reveals Model Adaptation Mechanism
Figure 2: Performance of LLaMA3.1-8B on CSR benchmarks as the percentage of selected modules varies for each PEFT method. Note the performance drops (e.g., DoRA at 60%) when redundant, high-sparsity modules are forced to update, indicating noise injection.
@@ 4.3 Analysis: Token Sparsity Reveals Model Adaptation Mechanism
Table 4: Performance of LLaMA3.1-8B on the commonsense reasoning task using various selection methods.
Sparsity Patterns Reflect Architecture Bias. In LLaMA3.1-8B, there are 32 layers in total. We compute the sparsity of q_proj, k_proj, and v_proj in each layer and take their average to obtain the sparsity distribution of TS-PEFT across modules (detailed results are provided in Appendix A.3). Overall, the sparsity of TS-LoRA and TS-AdaLoRA is relatively evenly distributed across layers, with standard deviations of 9.8 and 11.4, respectively. In contrast, TS-DoRA exhibits a much higher standard deviation of 41.3, showing a more “polarized” pattern. It should be noted that LoRA and AdaLoRA share the same weight structure despite having different rank allocation strategies, whereas DoRA adopts a distinct parameterization by decomposing weights into magnitude and direction. We hypothesize that this structural difference leads to the stronger selectivity observed in TS-DoRA at the module level, naturally raising the question: are the modules with lower sparsity more critical for downstream tasks?
@@
Redundant Updates Introduce Optimization Noise. To further validate that the identified high-sparsity modules are indeed redundant and potentially harmful, we conducted a cumulative selection analysis. We expanded the set of target modules in ascending order of sparsity (i.e., adding modules from “critical” to “redundant”). As illustrated in Figure 2, the conventional wisdom that ”more trainable parameters imply better performance” does not hold. For DoRA, performance peaks when only 50% of the modules are fine-tuned, while LoRA reaches its optimum at 80%. Remarkably, even utilizing just 20% of the modules yields results comparable to the optimal settings
Crucially, we observe distinct “performance cliffs” when redundant modules are forced to update. For instance, DoRA suffers a sharp performance drop of 6.6% when the target proportion increases from 50% to 60%. Similarly, LoRA experiences a 1.3% degradation when expanding from 90% to 100%. This phenomenon provides compelling evidence that updating modules characterized by high sparsity injects negative interference rather than meaningful adaptation. In contrast, AdaLoRA exhibits greater stability against these low-importance modules, likely due to its dynamic rank allocation mechanism which automatically limits the capacity assigned to redundant positions.
@@ 5 Conclusion
This paper revisits parameter-efficient fine-tuning from a token-level perspective and shows that standard PEFT methods exhibit substantial redundancy by uniformly updating all token positions within selected layers. We propose TS-PEFT, which introduces a learnable threshold gate to selectively apply PEFT updates to only about 40–60% of token positions, yet consistently matches or surpasses strong baselines. Our analyses further show that the resulting token-level sparsity patterns provide a meaningful signal of module importance, enabling more effective module selection under the same parameter budget and offering a theoretical basis for future hardware-software co-design targeting efficient sparse fine-tuning and inference.